OpenStudy (anonymous):

I have (-x^2 - p^2)e^xt*sin(pt+y) + (x^2 + p^2)x where x,p and y are constants, and t is the variable. Is there anyway to simplify this and make it equate to 0??

hartnn (hartnn):

(e^x)t or e^(xt)

hartnn (hartnn):

x,p,y are constants, so the form is Ae^(xt) sin (pt+y) +C = 0 why u want it = 0 ?

hartnn (hartnn):

X^2+p^2 gets cancelled out

OpenStudy (anonymous):

ok, this question probably seems a little incomplete, so i'll put the whole thing. a(t) = e^(-xt)*sin(pt+y), where x, p and y are constants. Show that \[\frac{ d^2a }{ dt^2 } + 2x \frac{ da }{ dt } + (x^2+p^2)a = \]

OpenStudy (anonymous):

AAAAAAAND i just realised where I went wrong.

hartnn (hartnn):

great! just my presence is enough :P

OpenStudy (anonymous):

that last x in the question was supposed to be an a, and then it ALL makes sense :P I'm an idiot.

hartnn (hartnn):

yes, it does :)